Adam is riding forward on electric scooter at 10m/s. He throws a ball to Robert , who is stationary. The ball travels at a speed of 8 m/s relative to Adam. How fast is the ball moving relative to Robert as he tries to catch it?
1
Expert's answer
2011-10-21T09:38:20-0400
Velocity is a vector value. So, speed is depends of Robert's position.For example,if he stays ahead to Adam:
v=v1+v2=10+8=18m/s
if he stays behind to Adam:
v=v1-v2=10-8=2m/s